# Christoph Bandt

### v1

@@ Abstract
Finite automata were used to determine multiple addresses in number systems and to find topological properties of self-affine tiles and finite type fractals. We join these two lines of research by axiomatically defining automata which generate topological spaces. Simple examples show the potential of the concept. Spaces generated by automata are topologically self-similar. Two basic algorithms are outlined. The first one determines automata for all $k$-tuples of equivalent addresses from the automaton for double addresses. The second one constructs finite topological spaces which approximate the generated space. Finally, we discuss the realization of automata-generated spaces as self-similar sets.
@@ 1 Overview
The basic principle of number systems is to assign addresses to points. We have an address map $\varphi:S\to X$ from a symbolic space $S$ to a set $X$ of points. The simplest symbolic space is the space of sequences $s=i_{1}i_{2}...$ from the digit set $D=\{0,1,...,m-1\}.$ Two symbol sequences $s,t$ are called equivalent if they address the same point: $\varphi(s)=\varphi(t).$ This paper studies algorithms which describe multiple addresses. If the set of pairs $(s,t)$ of equivalent addresses is generated by a finite automaton, the address map $\varphi$ is called automatic, and the topological quotient space $X=S/\varphi$ is automata-generated.
Gilbert [19, 20] started work in this direction in 1982. He constructed automata which determine double and triple expansions in number systems with complex bases $-n+i,n=1,2,...$ Thurston emphasized the connection between number systems, self-similar tilings and automata in his influential lecture notes [46]. This work was extended in numerous papers on canonical number systems (cf. [2] and [18, Section 2.4]) and on the topology of self-affine tiles in the plane [1, 13, 28, 29, 30, 31, 42] and in space [47]. Roughly speaking, self-affine tiles are the unit intervals of canonical number systems. Their vertices, which are interesting from a geometrical viewpoint, have three or more addresses.
Symbolic dynamics is a related field where addressing plays a central part. Milnor and Thurston [38] characterized unimodal maps on the interval by their kneading sequence which represents the double address of the critical point. There are special parameters where this address is preperiodic. For complex rational functions, Thurston discussed the symbolic description of Julia sets in the post-critically finite (p.c.f.) case [45]. They were considered as topologically self-similar sets by Kameyama [24]. In fractal geometry, p.c.f. sets have been frequently studied because of their simple structure. Brownian motion and differential equations can be defined on such spaces [25, 26, 43, 44]. All p.c.f. fractals are generated by automata, as shown in [39] and in Section 5.
@@
However, self-similar sets generated by similitudes are mainly studied in the plane. In three dimensions, similitudes seem to special. Several authors used a combinatorial approach to study fractal topology, putting aside metric features like similarity and Hausdorff dimension [5, 22, 24, 44]. Zhu and Rao [50] defined topology automata for fractal squares. We follow this line of research.
The new point is that we do not derive the automaton from a number system, tile or fractal. We take the automaton as starting point: as a tool to construct a topological space. In Section 2 we axiomatically introduce the class of automata which accept the double addresses in an address map. We give many simple examples and illustrations in Sections 3 and 5. Graphs were drawn by MATLAB, fractals by IFStile [37]. In Section 4 we prove that automata-generated spaces are topologically self-similar, even in a more general setting.
Two basic algorithms are described. In Section 6 we derive automata for the triple and $k$-tuple addresses from the original automaton for double addresses. For an automaton with three states and three digits, we find all points in $X$ with 4, 6 and 12 addresses. In Section 7 we construct the topological space $X$ from finite space approximations, unfolding the edge structure of the automaton. Since the automaton is finite, topological properties of the limit should be computable on a certain finite level. This is work in progress. In the final Section 9 we discuss the realization of an automata-generated space as a self-similar set in the complex plane. Many open questions appear in the last part of the paper. A long-term target is to establish databases of recursively defined topological objects, maintained and developed by computer.
This paper is dedicated to Christiane Frougny. Without her, it would not have been written. Many years ago, she invited me to Paris and introduced me to her numeration community which profoundly influenced my direction of research. I am still very grateful to Christiane. The automata in this paper are very similar to those studied by Frougny and Sakarovitch [18], but they serve a more geometric purpose.
@@ 2 The concept of a topology-generating automaton
Throughout, $D=\{0,1,...,m-1\}$ is our alphabet, the set of digits for numeration, denoted by $i,j,i_{k},u_{k},s_{k}.$ Words $i_{1}...i_{n}\in D^{n}$ will be denoted by $u,v,w$ and sequences $i_{1}i_{2}...$ by $s,t,$ or $\overline{u}=uuu...$ for periodic sequences. The set of words is $D^{*}=\bigcup_{n=0}^{\infty}D^{n}.$ The set of sequences is our symbolic space $S,$ the full one-sided shift over $D$ with the usual product topology. Elements of $S$ are called addresses.
$S=D^{\mathbb{N}}=\{s=s_{1}s_{2}...\,|\,s_{k}\in D\}\ .$
@@
Example 1 (Binary numbers)
Let $D=\{0,1\},X=[0,1]$ and $\varphi(i_{1}i_{2}...)=\sum_{k=1}^{\infty}i_{k}\cdot 2^{-k}.$ Then $\varphi(s)=\varphi(t)$ if and only if either $s=t$ or there is an integer $n\geq 0$ and $i_{1},...,i_{n}\in D$ such that $s=i_{1}...i_{n}0\overline{1}$ and $t=i_{1}...i_{n}1\overline{0},$ or conversely. The corresponding automaton $G$ is drawn in Figure 1. The initial state is $o,$ and the input alphabet is the set $D^{2}$ of pairs of symbols. A pair of words or sequences $(u,v)$ is accepted if there is a directed path of edges with labels $(u_{1},v_{1}),(u_{2},v_{2}),...$ in $G$ which starts in $o.$ For $u=v$ this path will consist of loops from $o$ to itself. When $u$ is lexikographically smaller than $v,$ the edge with label $(0,1)$ will lead to the state $right,$ and all following labels must be $(1,0).$ In this way, the automaton defines the equivalence of binary addresses and thus the topology of the interval.
Before we formally define our automata, we note how the equivalence of addresses is expressed on the level of words of length $n.$ For a word $u=u_{1}...u_{n}$ the corresponding cylinder set in the symbolic space is defined as
$S_{u}=\{u_{1}u_{2}...u_{n}s_{n+1}s_{n+2}...\,|\,s_{k}\in D\}\subseteq S\ .$
Proposition 1
@@
$\varphi(s)=\varphi(t)\quad\mbox{ if and only if }\ \varphi(S_{s_{1}...s_{n}})\cap\varphi(S_{t_{1}...t_{n}})\not=\emptyset\ \mbox{ for }\ n=1,2,...$
Proof. If $x=\varphi(s)=\varphi(t)$ then $x$ belongs to all intersections on the right-hand side. On the other hand, $\bigcap_{n}S_{t_{1}...t_{n}}=\{t\},$ and the $M_{n}=\varphi(S_{s_{1}...s_{n}})\cap\varphi(S_{t_{1}...t_{n}})$ form a decreasing sequence of closed sets. If they are all nonempty, $M=\bigcap_{n}M_{n}$ is nonempty by compactness, and must coincide with $\{\varphi(t)\}$ as well as with $\{\varphi(s)\}.$Box
The sets $X_{t_{1}...t_{n}}=\varphi(S_{t_{1}...t_{n}})$ will be called the pieces of level $n$ for the address map $\varphi.$ The proposition says that two sequences address the same point if and only if the corresponding pieces of level $n=1,2,...$ intersect each other. In example 1, pieces of level $n$ are binary intervals $[x,x+2^{-n}]$ with $x=\varphi(t_{1}...t_{n}\overline{0}).$ For $n=1$ we have $X_{0}=[0,\frac{1}{2}]$ and $X_{1}=[\frac{1}{2},1]$ with one common endpoint. If we start with the edge with label $(i,j)=(0,1)$ then $X_{j}$ is on the right of $X_{i}$ which explains the name of the state in Figure 1. For $n=2,$ only $X_{i1}$ and $X_{j0}$ intersect, where the latter remains on the right. In this way, the automaton is constructed. A state can be interpreted as a relative position of the two pieces.
Definition 1 (Topology-generating automaton)
A topology-generating automaton $G=(V,D^{2},E,o)$ consists of a finite directed graph $(V,E)$ where vertices represent states and edges represent transitions labelled by the input alphabet $D^{2},$ and $o$ denotes the initial state. The graph can contain loops and multiple edges which can be drawn as single edges with multiple labels. The following properties are required.
@@
    To each state $c$ there is an ‘inverse’ state $\pi(c)$ such that for every edge between vertices $b$ and $c$ labelled with $(i,j),$ there is an edge labelled with $(j,i)$ between $\pi(b)$ and $\pi(c).$ The map $\pi:V\to V$ fulfils $\pi(\pi(c))=c$ and $\pi(o)=o.$
@@
We use the convention that an input $(i,j)$ is accepted at state $c\in V$ if there is an edge with label $(i,j)$ starting in $c.$ A pair of words $(i_{1}...i_{n},j_{1}...j_{n})$ or sequences $(i_{1}i_{2}...,j_{1}j_{2}...)$ is accepted if there is a directed path of edges starting in $o$ with labels $(i_{1},j_{1}),(i_{2},j_{2}),...$ The language of accepted pairs of words and the set of accepted pairs of sequences are denoted by $L(G)\subseteq(D^{2})^{*}$ and $L^{\infty}(G)\subseteq S^{2},$ respectively.
We used common terminology as in [12, 16]. While automata for number systems as treated by Frougny and Sakarovitch [18] use the input alphabet $D,$ we need $D^{2}=D\times D$ to process pairs of symbol sequences. The absence of a set of final states was explained by Thurston as follows: “The language $L(G)$ is prefix-closed if every prefix of a word in $L(G)$ is also in $L(G),$ or in other words, if every non-accept state has arrows only to other non-accept states. In such a case, we may collapse all non-accept states in a single fail state, with all arrows leading back to itself. It is convenient to omit the fail state and all roads leading to it. Whenever an input gives you directions where there is no corresponding arrow, you immediately fail with no chance for reinstatement.” [46, p. 31].
According to Proposition 1 we want to describe the relation $\varphi(S_{u})\cap\varphi(S_{v})\not=\emptyset$ for words $u,v$ on $D^{2}.$ This relation is prefix-closed since $S_{u^{\prime}}\supseteq S_{u}$ for any prefix $u^{\prime}$ of $u.$ So Thurston’s convention applies here. Moreover, if the relation is fulfilled for $(u,v)$ then it must also be fulfilled by $(ui,vj)$ for some $(i,j)\in D^{2},$ since $S_{u}=\bigcup_{i\in D}S_{ui}.$ This explains the outgoing edges in property 1. These edges guarantee that each directed path of edges can be extended indefinitely, in a finite graph through directed cycles. Each vertex should have an incoming edge, and should be reached from $o$ because otherwise it is obsolete.
@@
Given an automaton with these properties, we can define the quotient space $X$ as the set of all equivalence classes $x_{s}$ of addresses $s.$ The address map $\varphi$ will assign to each address $s\in S$ its equivalence class.
@@
One way to define the quotient topology on $X$ is to say that a sequence $(x_{s^{n}})_{n=1,2,...}$ with $s^{n}\in S$ converges to $x_{s}$ in $X$ if and only if $s^{n}$ converges to $s$ in $S.$ However, there can be $t$ with $x_{s}=x_{t}$ and $t^{n}$ with $x_{s^{n}}=x_{t^{n}}.$ We show that the convergence does not depend on the choice of representatives.
Proposition 2
@@
    Let $(s^{n})_{n=1,2,...}$ and $(t^{n})_{n=1,2,...}$ be convergent sequences of addresses with limit sequences $s$ and $t$ in $S.$ If $(s^{n},t^{n})$ is accepted by $G$ for $n=1,2,...$ then $(s,t)$ is also accepted by $G.$
@@
Proof. (i) In $S$ we have coordinatewise convergence. Thus the first $k$ digits of $s^{n}$ must agree with the digits $s_{1}s_{2}...s_{k}$ of the limit sequence $s$ for all $n$ greater than some $n(k).$ Similarly, $t^{n}$ starts with $t_{1}t_{2}...t_{k}$ for large enough $n.$ Since $(s^{n},t^{n})$ is accepted by $G$ for all $n,$ this implies $\varphi(S_{s_{1}...s_{k}})\cap\varphi(S_{t_{1}...t_{k}})\not=\emptyset.$ This holds for all $k,$ so $(s,t)$ is accepted by $G.$
(ii) By definition of the quotient topology, $\varphi$ is continuous. The space $S$ is compact, and a continuous image of a compact space is always compact. A space is Hausdorff if limits are uniquely determined. This is what we proved in (i) for the quotient topology of $X.$ Since $S$ has a countable base, this holds for $X$ and implies that $X$ is metrizable. $\Box$
@@
Proposition 3
Let all equivalence classes of addresses be finite, and let $(s,t)$ be a pair of different sequences accepted by the automaton $G.$ Then $s$ is non-periodic and cannot be written as $s=ut$ for a word $u\in D^{*}.$
Proof. If a periodic sequence $s=\overline{w}$ is identified with $t,$ then $s=ws$ is identified with $wt$ by property 4. By induction $s$ is also identified with $w^{k}t$ for $k=2,3,...\,.$ Thus the equivalence class of a periodic sequence is infinite.
If $s$ is identified with $us$ then by property 4 $us$ is identified with $u^{2}s$ and $u^{2}s$ with $u^{3}s$ and so on. Again we have an infinity of equivalent addresses $u^{k}s,k=0,1,2,...,$ and $\overline{u}$ will also belong to this class because the equivalence relation is closed according to Proposition 2. So we have the same case as above. $\Box$
A particular case appears when there is a path from $o$ to $o$ starting with labels $i_{1}\not=j_{1}.$ If $u,v$ are the words addressing this path, then not only $\overline{u}$ and $\overline{v}$ are identified. Every sequence in $\{u,v\}^{\infty}$ will belong to the equivalence class which therefore has the cardinality of the continuum. There are important number systems like $\beta$-numeration with golden mean base which have this property. They are also interesting from the topological viewpoint but require methods which are not studied here. For that reason, we assume that the loops of property 4 are the only incoming edges to $o.$

### v2

@@ Abstract
Finite automata were used to determine multiple addresses in number systems and to find topological properties of self-affine tiles and finite type fractals. We join these two lines of research by axiomatically defining automata that generate topological spaces. Simple examples show the potential of the concept. Spaces generated by automata are topologically self-similar. Two basic algorithms are outlined. The first one determines automata for all $k$-tuples of equivalent addresses from the automaton for double addresses. The second one constructs finite topological spaces which approximate the generated space. Finally, we discuss the realization of automata-generated spaces as self-similar sets.
keywords:
Self-similarity, finite-state automaton, topological quotient space, multiple addresses, symbolic dynamics.
[…]
@@ 1 Overview
The basic principle of number systems is to assign addresses to points. We have an address map $\varphi:S\to X$ from a symbolic space $S$ to a set $X$ of points. The simplest symbolic space is the space of sequences $s=i_{1}i_{2}\dots$ from the digit set $D=\{0,1,\dots,m-1\}$. Two symbol sequences $s,t$ are called equivalent if they address the same point: $\varphi(s)=\varphi(t)$. This paper studies algorithms that describe multiple addresses. If the set of pairs $(s,t)$ of equivalent addresses is generated by a finite automaton, the address map $\varphi$ is called automatic, and the topological quotient space $X=S/\varphi$ is automata-generated.
Gilbert [20, 21] started work in this direction in 1982. He constructed automata which determine double and triple expansions in number systems with complex bases $-n+i$, for $n=1,2,\dots$ Thurston emphasized the connection between number systems, self-similar tilings and automata in his influential lecture notes [48]. This work was extended in numerous papers on canonical number systems (cf. [2] and [19, Section 2.4]) and on the topology of self-affine tiles in the plane [1, 14, 30, 29, 31, 32, 44] and in space [49]. Roughly speaking, self-affine tiles are the unit intervals of canonical number systems. Their vertices, which are interesting from a geometrical viewpoint, have three or more addresses.
Symbolic dynamics is a related field where addressing plays a central part. Milnor and Thurston [39] characterized unimodal maps on the interval by their kneading sequence which represents the double address of the critical point. There are special parameters where this address is preperiodic. For complex rational functions, Thurston discussed the symbolic description of Julia sets in the post-critically finite (p.c.f.) case [47]. They were considered topologically self-similar sets by Kameyama [25]. In fractal geometry, p.c.f. sets have been frequently studied because of their simple structure. Brownian motion and differential equations can be defined on such spaces [26, 27, 45, 46]. All p.c.f. fractals are generated by automata, as shown in [41] and in Section 5.
@@
However, self-similar sets generated by similitudes are mainly studied in the plane. In three dimensions, similitudes seem to be too special. Several authors used a combinatorial approach to study fractal topology, putting aside metric features like similarity and Hausdorff dimension [3, 7, 23, 25, 40, 46]. Zhu and Rao [52] defined topology automata for fractal squares. We follow this line of research.
The new point is that we do not derive the automaton from a number system, tile or fractal. We take the automaton as a starting point: as a tool to construct a topological space. In Section 2 we axiomatically introduce the class of automata which accepts the double addresses in an address map. We give many simple examples and illustrations in Sections 3 and 5. Graphs were drawn by MATLAB, fractals by IFStile [38]. In Section 4 we prove that automata-generated spaces are topologically self-similar, even in a more general setting.
Two basic algorithms are described. In Section 6 we derive automata for the triple and $k$-tuple addresses from the original automaton for double addresses. For an automaton with three states and three digits, we find all points in $X$ with 4, 6, and 12 addresses. In Section 7 we construct the topological space $X$ from finite space approximations, unfolding the edge structure of the automaton. Since the automaton is finite, topological properties of the limit should be computable on a certain finite level. This is a work in progress. In the final Section 9 we discuss the realization of an automata-generated space as a self-similar set in the complex plane. Many open questions appear in the last part of the paper. A long-term target is to establish databases of recursively defined topological objects, maintained and developed by computer.
The automata discussed in this paper are quite similar to those studied by Frougny and Sakarovitch [19], but they are used for a more geometric purpose.
@@ 2 The concept of a topology-generating automaton
Throughout, $D=\{0,1,\dots,m-1\}$ is our alphabet, the set of digits for numeration, denoted by $i,j,i_{k},u_{k},s_{k}$. Words $i_{1}\dots i_{n}\in D^{n}$ will be denoted by $u,v,w$ and sequences $i_{1}i_{2}\dots$ by $s,t$, or $\overline{u}=uuu\dots$ for periodic sequences. The set of words is $D^{*}=\bigcup_{n=0}^{\infty}D^{n}$. The set of sequences is our symbolic space $S$, the full one-sided shift over $D$ with the usual product topology. Elements of $S$ are called addresses.
$S=D^{\mathbb{N}}=\{s=s_{1}s_{2}\dots\,|\,s_{k}\in D\}.$
@@
Example 2.1 (Binary numbers).
Let $D=\{0,1\},X=[0,1]$ and $\varphi(i_{1}i_{2}\dots)=\sum_{k=1}^{\infty}i_{k}\cdot 2^{-k}$. Then $\varphi(s)=\varphi(t)$ if and only if either $s=t$ or there is an integer $n\geq 0$ and $i_{1},\dots,i_{n}\in D$ such that $s=i_{1}\dots i_{n}0\overline{1}$ and $t=i_{1}\dots i_{n}1\overline{0}$, or conversely. The corresponding automaton $G$ is drawn in Figure 1. The initial state is $o$, and the input alphabet is the set $D^{2}$ of pairs of symbols. A pair of words or sequences $(u,v)$ is accepted if there is a directed path of edges with labels $(u_{1},v_{1}),(u_{2},v_{2}),\dots$ in $G$ which starts in $o$. For $u=v$ this path will consist of loops from $o$ to itself. When $u$ is lexicographically smaller than $v$, the edge with label $(0,1)$ will lead to the state right, and all following labels must be $(1,0)$. In this way, the automaton defines the equivalence of binary addresses and thus the topology of the interval.
Before we formally define our automata, we note how the equivalence of addresses is expressed on the level of words of length $n$. For a word $u=u_{1}\dots u_{n}$ the corresponding cylinder set in the symbolic space is defined as
$S_{u}=\{u_{1}u_{2}\dots u_{n}s_{n+1}s_{n+2}\dots\,|\,s_{k}\in D\}\subseteq S.$
Proposition 2.2.
@@
$\varphi(s)=\varphi(t)\mbox{ if and only if }\ \varphi(S_{s_{1}\dots s_{n}})\cap\varphi(S_{t_{1}\dots t_{n}})\not=\emptyset\ \mbox{ for }\ n=1,2,\dots$
Proof 2.3.
If $x=\varphi(s)=\varphi(t)$ then $x$ belongs to all intersections on the right-hand side. On the other hand, $\bigcap_{n}S_{t_{1}\dots t_{n}}=\{t\}$, and the $M_{n}=\varphi(S_{s_{1}\dots s_{n}})\cap\varphi(S_{t_{1}\dots t_{n}})$ form a decreasing sequence of closed sets. If they are all nonempty, $M=\bigcap_{n}M_{n}$ is nonempty by compactness, and must coincide with $\{\varphi(t)\}$ as well as with $\{\varphi(s)\}$.
The sets $X_{t_{1}\dots t_{n}}=\varphi(S_{t_{1}\dots t_{n}})$ will be called the pieces of level $n$ for the address map $\varphi$. The proposition says that two sequences address the same point if and only if the corresponding pieces of level $n=1,2,\dots$ intersect each other. In example 2.1, pieces of level $n$ are binary intervals $[x,x+2^{-n}]$ with $x=\varphi(t_{1}\dots t_{n}\overline{0})$. For $n=1$ we have $X_{0}=[0,\frac{1}{2}]$ and $X_{1}=[\frac{1}{2},1]$ with one common endpoint. If we start with the edge with label $(i,j)=(0,1)$ then $X_{j}$ is on the right of $X_{i}$ which explains the name of the state in Figure 1. For $n=2$, only $X_{i1}$ and $X_{j0}$ intersect, where the latter remains on the right. In this way, the automaton is constructed. A state can be interpreted as a relative position of the two pieces.
Definition 2.4 (Topology-generating automaton).
A topology-generating automaton, denoted by $G=(V,D^{2},E,o)$, consists of a finite directed graph $(V,E)$ where vertices represent states and edges represent transitions labelled by the input alphabet $D^{2}$, and $o$ denotes the initial state. The graph can contain loops and multiple edges which can be drawn as single edges with multiple labels. The following properties are required.
@@
    To each state $c$ there is an ‘inverse’ state $c-$ such that for every edge between vertices $b$ and $c$ labelled with $(i,j)$, there is an edge labelled with $(j,i)$ between $b-$ and ${c-}$. The conditions $o-=o$ and $(c-)-=c$ for $c\in V$ are fulfilled.
@@
We use the convention that an input $(i,j)$ is accepted at state $c\in V$ if there is an edge with label $(i,j)$ starting in $c$. A pair of words $(i_{1}\dots i_{n},j_{1}\dots j_{n})$ or sequences $(i_{1}i_{2}\dots,j_{1}j_{2}\dots)$ is accepted if there is a directed path of edges starting in $o$ with labels $(i_{1},j_{1}),(i_{2},j_{2}),\dots$. The language of accepted pairs of words and the set of accepted pairs of sequences are denoted by $L(G)\subseteq(D^{2})^{*}$ and $L^{\infty}(G)\subseteq S^{2}$, respectively.
We used common terminology as in [13, 17]. While automata for number systems as treated by Frougny and Sakarovitch [19] use the input alphabet $D$, we need $D^{2}=D\times D$ to process pairs of symbol sequences. The absence of a set of final states was explained by Thurston as follows: “The language $L(G)$ is prefix-closed if every prefix of a word in $L(G)$ is also in $L(G)$, or in other words, if every non-accept state has arrows only to other non-accept states. In such a case, we may collapse all non-accept states in a single fail state, with all arrows leading back to itself. It is convenient to omit the fail state and all roads leading to it. Whenever an input gives you directions where there is no corresponding arrow, you immediately fail with no chance for reinstatement.” [48, p. 31].
According to Proposition 2.2 we want to describe the relation $\varphi(S_{u})\cap\varphi(S_{v})\not=\emptyset$ for words $u,v$ on $D^{2}$. This relation is prefix-closed since $S_{u^{\prime}}\supseteq S_{u}$ for any prefix $u^{\prime}$ of $u$. So Thurston’s convention applies here. Moreover, if the relation is fulfilled for $(u,v)$ then it must also be fulfilled by $(ui,vj)$ for some $(i,j)\in D^{2}$, since $S_{u}=\bigcup_{i\in D}S_{ui}$. This explains the outgoing edges in property 1. These edges guarantee that each directed path of edges can be extended indefinitely, in a finite graph through directed cycles. Each vertex should have an incoming edge and should be reached from $o$ because otherwise it is obsolete.
@@
Given an automaton with these properties, we formally define the quotient space $X$ as the set of all equivalence classes $x_{s}$ of addresses $s$. The address map $\varphi$ will assign to each address $s\in S$ its equivalence class.
@@
One way to define the quotient topology on $X$ is to say that a sequence $(x_{s^{n}})_{n=1,2,\dots}$ with $s^{n}\in S$ converges to $x_{s}$ in $X$ if and only if $s^{n}$ converges to $s$ in $S$. However, there can be $t$ with $x_{s}=x_{t}$ and $t^{n}$ with $x_{s^{n}}=x_{t^{n}}$. We show that the convergence does not depend on the choice of representatives.
Proposition 2.5.
@@
    Let $(s^{n})_{n=1,2,\dots}$ and $(t^{n})_{n=1,2,\dots}$ be convergent sequences of addresses with limit sequences $s$ and $t$ in $S$. If $(s^{n},t^{n})$ is accepted by $G$ for $n=1,2,\dots$. then $(s,t)$ is also accepted by $G$.
@@
Proof 2.6.
  1. (i)
    In $S$ we have coordinatewise convergence. Thus the first $k$ digits of $s^{n}$ must agree with the digits $s_{1}s_{2}\cdots s_{k}$ of the limit sequence $s$ for all $n$ greater than some $n(k)$. Similarly, $t^{n}$ starts with $t_{1}t_{2}\cdots t_{k}$ for large enough $n$. Since $(s^{n},t^{n})$ is accepted by $G$ for all $n$, this implies $\varphi(S_{s_{1}\cdots s_{k}})\cap\varphi(S_{t_{1}\cdots t_{k}})\not=\emptyset$. This holds for all $k$, so $(s,t)$ is accepted by $G$.
  2. (ii)
    By definition of the quotient topology, $\varphi$ is continuous. The space $S$ is compact, and a continuous image of a compact space is always compact. A space is Hausdorff if limits are uniquely determined. This is what we proved in (i) for the quotient topology of $X$. Since $S$ has a countable base, this holds for $X$ and implies that $X$ is metrizable.
@@
Proposition 2.7.
Let all equivalence classes of addresses be finite, and let $(s,t)$ be a pair of different sequences accepted by the automaton $G$. Then $s$ is not a periodic sequence and cannot be written as $s=ut$ for a word $u\in D^{*}$.
Proof 2.8.
If a periodic sequence $s=\overline{w}$ is identified with $t$, then $s=ws$ is identified with $wt$ by property 4. By induction $s$ is also identified with $w^{k}t$ for $k=2,3,\dots\,$. Thus the resulting equivalence class is infinite.
If $t$ is identified with $ut$ then by property 4 $ut$ is identified with $u^{2}t$ and $u^{2}t$ with $u^{3}t$ and so on. Again we have an infinity of equivalent addresses $u^{k}t,k=0,1,2,\dots$, and $\overline{u}$ will also belong to this class because the equivalence relation is closed according to Proposition 2.5. So we have the same conclusion as above.
A particular case appears when there is a path from $o$ to $o$ starting with labels $i_{1}\not=j_{1}$. If $u,v$ are the words addressing this path, then not only $\overline{u}$ and $\overline{v}$ are identified. Every sequence in $\{u,v\}^{\infty}$ will belong to the equivalence class which therefore has the cardinality of the continuum. There are important number systems like $\beta$-numeration with a golden mean base which have this property. They are also interesting from the topological viewpoint but require methods that are not studied here. For that reason, we assume that the loops of property 4 are the only incoming edges to $o$.